OpenStudy (anonymous):

Which is the exact value of the expression sqrt{32} - sqrt{50} + sqrt{128}? A.sqrt[2]{7} B.sqrt[7]{2} C.sqrt[22]{2} D.sqrt[2]{55}

OpenStudy (anonymous):

\[\sqrt{32} - \sqrt{50} + \sqrt{128}\]

OpenStudy (primeralph):